Let f be a function from Z to R, such that f(x)=x/10, then f is

a) an increasing function   

b) a strictly increasing function

c) a decreasing function

d) an onto function


1
Expert's answer
2021-05-07T14:21:50-0400

Let "f" be a function from "\mathbb Z" to "\mathbb R", such that "f(x)=\frac{x}{10}". If "x<y", then "\frac{x}{10}<\frac{y}{10}", and hence "f(x)<f(y)". It follows that "f" is a strictly increasing function.


The function "f" is not an onto function. Indeed, for "y=\frac{1}{20}" the equation "f(x)=\frac{1}{20}," which is equivalent to "\frac{x}{10}=\frac{1}{20}" and hence to "x=\frac{1}{2}," has no solution in the set "\mathbb Z" of integer numbers. Therefore, there is no integer number "x" such that "f(x)=\frac{1}{20}."